# Less Is More for Multi-Step Logical Reasoning of LLM Generalisation Under Rule Removal, Paraphrasing, and Compression

### v1

@@ Abstract
Large language models (LLMs) excel across many natural language tasks, yet their generalisation to structural perturbations in logical contexts remains poorly understood. We introduce a controlled evaluation framework that probes reasoning reliability through four targeted stress tests: (1) rule deletion, removing either redundant or essential rules from a multi-step inference chain; (2) contradictory evidence injection; (3) logic-preserving rewrites generated through several families of equivalence laws (contrapositive, double negation, implication, De Morgan, identity, and commutativity); and (4) multi-law equivalence stacking that introduces 2–5 simultaneous logical transformations.
Across three representative model families—BERT, Qwen2, and LLaMA-like models—our experiments reveal a strikingly consistent pattern: all models achieve perfect accuracy on the base tasks and remain fully generalise to redundant rule deletion and all equivalence-based rewrites (single or multi-law), but fail sharply under essential rule deletion (dropping to 25% accuracy) and collapse completely in the presence of explicit contradictions (0% accuracy). These results demonstrate that LLMs possess stable invariance to semantic-preserving logical transformations, yet remain fundamentally brittle to missing or conflicting evidence. Our framework provides a clean diagnostic tool for isolating such reasoning failure modes and highlights persistent gaps in the logical generalisation abilities of current LLMs.
@@ Introduction
Reasoning over structured rule systems is a fundamental capability required for reliable decision-making and language understanding. Although large language models (LLMs) achieve impressive performance across numerous natural language benchmarks, it is still unclear to what extent these models perform true logical reasoning rather than relying on memorized patterns or superficial correlations. Much prior work evaluates LLMs in settings where all relevant information is present in canonical form; however, real-world reasoning often involves incomplete information, rephrased or transformed rules, redundant evidence, and even explicit contradictions. Understanding how LLMs behave under such perturbations is essential for both theoretical insight and safe deployment.
@@
To address this gap, we introduce a highly controlled evaluation framework that isolates core aspects of rule-based reasoning. The framework is built around a canonical multi-step logical chain (e.g., green → cold → rough → young → nice) that requires models to integrate several rules in sequence. From this base structure, we generate a family of targeted variants that probe different dimensions of reasoning generalisation:
Rule deletion. We remove either redundant rules—whose absence should not change conclusions—or essential rules that constitute the backbone of the inference chain. This tests whether models distinguish between logically important and unimportant components.
Contradictory evidence injection. We introduce explicit facts that conflict with conclusions implied by the rule chain, requiring the model to reconcile or prioritize between mutually inconsistent pieces of information.
@@
Logic-preserving transformations. We rewrite rules using several major families of logical equivalence laws, including contraposition, double negation, implication-to-disjunction, De Morgan’s laws, identity, and commutativity of disjunction. These variants preserve the underlying semantics while changing surface form, allowing us to test whether LLMs track logical content rather than specific linguistic templates.
@@
Multi-law equivalence stacking. We introduce conditions in which 2–5 logical equivalence laws are applied simultaneously to create highly redundant, cluttered, and structurally complex rule sets, all of which remain logically equivalent to the base rules.
@@
We evaluate this framework across three distinct model families—BERT, Qwen2, and LLaMA-like models—using identical training and testing pipelines. Despite architectural differences, all models exhibit the same qualitative and quantitative behavior:
  • Perfect reasoning on the base tasks (100% accuracy)
  • Full generalisation to redundant rule deletion (100% accuracy)
  • Catastrophic failure under essential rule deletion (25% accuracy)
  • Complete collapse under explicit contradictions (0% accuracy)
  • Full invariance to all six classes of logical equivalence rewrites (100% accuracy)
  • Full invariance even under multi-law equivalence stacking (100% accuracy)
Together, these results reveal a consistent pattern: LLMs correctly identify which rules are structurally necessary, generalise flawlessly across a wide spectrum of logically equivalent transformations, but lack the capacity to adjust reasoning when evidence is incomplete or contradictory. Rather than flexibly integrating conflicting information, models tend to complete the rule chain whenever possible, even when doing so contradicts stated facts.
Our contributions provide a clean and extensible diagnostic tool for probing reasoning behavior, enabling precise identification of reasoning failure modes. The findings highlight a persistent gap between surface-level generalisation to paraphrase-like variation and broader generalisation to structural perturbation or conflict, underscoring the need for models and training strategies that can reason reliably under minimal, missing, or inconsistent evidence.
@@ 4. Evaluation & Analysis
base 1.00 0.00 1.00 0.00 1.00 0.00
variant1 1.00 0.00 1.00 0.00 1.00 0.00
variant2 0.25 -0.75 0.25 -0.75 0.25 -0.75
variant3 0.00 -1.00 0.00 -1.00 0.00 -1.00
variant4-contrapositive 1.00 0.00 1.00 0.00 1.00 0.00
variant4-double-negation 1.00 0.00 1.00 0.00 1.00 0.00
variant4-implication 1.00 0.00 1.00 0.00 1.00 0.00
variant4-de-morgan 1.00 0.00 1.00 0.00 1.00 0.00
variant4-identity 1.00 0.00 1.00 0.00 1.00 0.00
variant4-commutativity 1.00 0.00 1.00 0.00 1.00 0.00
variant4-multi (2–5 laws) 1.00 0.00 1.00 0.00 1.00 0.00
@@
Table 1 presents a unified comparison of BERT, Qwen2, and TinyLlama across all structural variants in our benchmark. Despite architectural differences, all three models exhibit an almost perfectly aligned pattern of behavior. Each model achieves 100% accuracy on the base reasoning task and retains this performance under redundant rule removal (Variant 1) as well as all logic-preserving equivalence transformations, including six single-law rewrites and multi-law (2–5) equivalence stacking. These conditions yield $\Delta=0$, confirming complete invariance to semantic-preserving logical reformulations.
In contrast, performance collapses when essential rule information is removed (Variant 2), with accuracy dropping uniformly to 25% for all models, corresponding to a $\Delta$ of 0.75-0.75. Even more strikingly, all models fail entirely under contradictory facts (Variant 3), scoring 0% accuracy and yielding the largest degradation from baseline ($\Delta=-1.00$). This cross-model consistency indicates that generalisation to logical equivalence and brittleness to structural disruptions are not tied to model size or architecture, but instead reflect systematic properties of current LLM reasoning mechanisms. The variations in our framework reveal key properties of LLM-based reasoning. By systematically manipulating the logical environment—removing rules, introducing contradictions, and applying logic-preserving equivalence transformations—we expose dimensions of generalisation and brittleness often hidden in more complex benchmarks. Together, these findings clarify what contemporary LLMs can and cannot generalise in rule-based reasoning.
@@
The progression from identical conclusions (Variation 1) to partially disrupted conclusions (Variation 2) to complete breakdown (Variation 3) reveals the fine-grained sensitivity of LLM reasoning to structural properties of the rule set. In Variation 1, the removal of a redundant rule leaves all models completely unaffected—indicating that LLMs can identify when a rule does not contribute new inferential content. In sharp contrast, Variation 2 demonstrates that eliminating a single essential step in the reasoning chain causes accuracy to collapse to chance levels, suggesting that LLMs cannot reconstruct missing inferential links even when the rest of the context is intact.
Variation 3 highlights a different form of fragility: all models overwhelmingly prefer rule-driven inferences over explicitly stated contradictory facts. This behavior implies that LLMs tend to follow the internal logic of the given rule system rather than adjusting their conclusions in response to new or conflicting information—an important consideration for applications involving dynamic or noisy environments.
The invariance observed across all logic-preserving rewrites—whether single-law or multi-law—shows that LLMs possess semantic stability, treating contraposition, double negation, implication re-expression, De Morgan transformations, and other equivalence laws as interchangeable. This indicates that they capture structural relationships between logically equivalent formulations, even though they falter with incomplete or inconsistent evidence. Overall, LLMs remain generalise to semantic-preserving transformations but extremely brittle to structural disruptions, a duality with important implications for AI safety: systems that appear reliable under surface-level variation may still fail when essential information is missing or contradictory, highlighting the need for methods that compensate for incomplete reasoning or reconcile conflicting evidence.
@@ Conclusion
This work introduced a controlled framework for evaluating LLM generalisation under systematic perturbations to logical reasoning. By constructing multi-step inference chains and generating variants involving rule deletion, contradiction injection, and multiple families of logic-preserving transformations, we isolated reasoning behaviors often obscured in prior benchmarks. Across BERT, Qwen2, and LLaMA-like models, we observed perfect performance and complete invariance to semantic-preserving rewrites, but sharp failures when essential rules were removed and total collapse under explicit contradictions. These results highlight a clear gap between surface-level generalisation and deeper logical competence, underscoring the need for models that can integrate incomplete information, handle conflicting evidence, and recover missing reasoning steps. Our framework provides a foundation for improving both LLM reasoning reliability and evaluation methodology.
@@ Appendix A Appendix
    • Rule 4: If someone is not young then they are not rough. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$

### v2

@@ Abstract
Large language models (LLMs) achieve strong performance on many natural language tasks, yet their generalisation under structured perturbations of logical rule systems remains insufficiently characterised. We present a controlled evaluation framework that probes reasoning reliability through four stress tests: (1) rule deletion, removing redundant versus essential rules from a multi-step inference chain; (2) contradictory evidence injection; (3) logic-preserving rewrites based on equivalence laws (contraposition, double negation, implication-to-disjunction, De Morgan, identity, and commutativity); and (4) multi-law equivalence stacking that composes 2–5 transformations.
Across three representative model families—BERT, Qwen2, and LLaMA-like models—all models attain Acc=1.0000=1.0000 on the base split and show no degradation under redundant rule deletion. In contrast, essential rule deletion yields a pronounced decrease to near-chance performance, and injecting explicit contradictions reduces accuracy to 0.0000. Under logic-preserving rewrites, accuracy is largely preserved for single-law transformations with only small degradations in a few cases, whereas multi-law stacking exposes model-dependent sensitivity: BERT matches the base condition, TinyLlama shows only marginal degradation, and Qwen2 exhibits a substantial drop.
Overall, the results indicate that contemporary LLMs are generally stable under semantic-preserving reformulations, yet remain brittle to missing or inconsistent evidence and may degrade under composed logical transformations depending on the model family. The proposed framework provides a concise diagnostic tool for isolating these failure modes and for evaluating logical generalisation beyond surface-form variation.
@@ Introduction
Reasoning over structured rule systems is a foundational capability for reliable decision-making and language understanding. Although large language models (LLMs) perform strongly on many benchmarks, it remains unclear to what extent such performance reflects robust logical inference as opposed to distributional heuristics. Moreover, practical reasoning settings frequently involve incomplete information, redundant premises, rephrased or logically transformed rules, and even explicit contradictions. Characterising model behaviour under these controlled perturbations is therefore important for both scientific understanding and safe deployment.
@@
To address this gap, we introduce a controlled evaluation framework that isolates key aspects of rule-based reasoning. The benchmark is built around a canonical multi-step inference chain (e.g., green $\rightarrow$ cold $\rightarrow$ rough $\rightarrow$ young $\rightarrow$ nice) requiring multi-hop integration of rules. From each base instance, we generate targeted variants designed to probe distinct dimensions of generalisation:
Rule deletion. We remove either redundant rules (whose removal should preserve derivability) or essential rules required to complete the inference chain, thereby testing sensitivity to missing inferential links.
Contradictory evidence injection. We inject explicit facts that conflict with rule-derived conclusions, testing whether models can accommodate inconsistency rather than defaulting to a single inference path.
@@
Logic-preserving transformations. We rewrite rules using standard equivalence laws (contraposition, double negation, implication-to-disjunction, De Morgan, identity, and commutativity). These variants preserve semantics while altering surface form, allowing us to test whether models track logical content beyond templated phrasing.
@@
Multi-law equivalence stacking. We compose 2–5 equivalence laws simultaneously to create structurally more complex but logically equivalent rule sets, testing robustness under composed transformations.
@@
We evaluate the framework on three model families—BERT, Qwen2, and LLaMA-like models—under identical training and evaluation pipelines. Despite architectural differences, the observed behaviour is broadly consistent:
  • Base reasoning is solved at Acc=1.0000=1.0000.
  • Redundant rule deletion yields no measurable degradation.
  • Essential rule deletion produces a marked decrease to near-chance performance.
  • Injecting explicit contradictions reduces accuracy to 0.0000.
  • Single-law equivalence rewrites are handled robustly, with only minor degradations in a small number of cases.
  • Multi-law stacking exposes model-dependent sensitivity: some models retain baseline accuracy, whereas others exhibit substantial degradation.
Together, these results highlight an asymmetry: contemporary LLMs are generally stable under semantic-preserving rewrites of rules, yet highly sensitive to missing or inconsistent evidence, and may degrade under composed logical transformations. The proposed framework provides an extensible, fine-grained diagnostic suite for isolating these failure modes, complementing existing reasoning benchmarks that often conflate multiple sources of variation.
@@ 4. Evaluation & Analysis
base 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant1 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant2 0.2950 -0.7050 0.2500 -0.7500 0.2500 -0.7500
variant3 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
variant4-contrapositive 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-double-negation 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-implication 1.0000 0.0000 0.9525 -0.0475 1.0000 0.0000
variant4-de-morgan 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-identity 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-commutativity 0.9925 -0.0075 1.0000 0.0000 1.0000 0.0000
variant4-multi (2–5 laws) 1.0000 0.0000 0.6450 -0.3550 0.9925 -0.0075
@@
Table 1 reports accuracy (Acc) and deviation from the base condition ($\Delta=\mathrm{Acc}_{\mathrm{variant}}-\mathrm{Acc}_{\mathrm{base}}$) for BERT, Qwen2, and TinyLlama across all structural variants. All models achieve Acc=1.0000=1.0000 on the base split and exhibit no degradation under redundant rule removal (Variant 1; $\Delta=0$). By contrast, removing an essential rule (Variant 2) yields a substantial drop (BERT: 0.2950, $\Delta=-0.7050$; Qwen2/TinyLlama: 0.2500, $\Delta=-0.7500$), indicating strong sensitivity to missing inferential links. Injecting explicit contradictory facts (Variant 3) reduces accuracy to 0.0000 for all models ($\Delta=-1.0000$), suggesting that the models do not reliably revise conclusions in the presence of inconsistency.
For logic-preserving single-law rewrites (Variant 4), performance is largely preserved: most conditions remain at 1.0000 accuracy, with only small degradations in a few settings (e.g., BERT under commutativity: 0.9925, $\Delta=-0.0075$; Qwen2 under implication-to-disjunction: 0.9525, $\Delta=-0.0475$). Under multi-law equivalence stacking (2–5 laws), robustness becomes model-dependent: BERT matches the base condition (1.0000), TinyLlama shows only a marginal reduction (0.9925, $\Delta=-0.0075$), whereas Qwen2 exhibits a pronounced decrease (0.6450, $\Delta=-0.3550$). This contrast suggests that compositional logical rewrites can introduce effective difficulty beyond single transformations, even when semantic equivalence is preserved.
@@
Variants 1–3 isolate model sensitivity to perturbations that preserve versus disrupt derivability. In Variant 1 (redundant rule deletion), all models match the base accuracy, consistent with the expectation that removing unused rules should not affect entailment. In Variant 2 (essential rule deletion), accuracy drops to near-chance levels, suggesting limited premise tracking: models often behave as if the missing link were still present and tend to “complete” the original inference pattern even when intermediate support is removed. This behaviour is consistent with reliance on learned chain templates rather than verifying support for each queried conclusion. In Variant 3 (contradictory evidence), all models score 0.0000, indicating that they do not reliably adopt contradiction-aware conservative reasoning under the benchmark’s scoring semantics.
For logic-preserving rewrites (Variant 4), performance is largely preserved for single-law transformations, with only minor degradations in a few cases (e.g., implication-to-disjunction or commutativity). However, multi-law stacking exposes model-dependent robustness: BERT remains at the base level, TinyLlama shows only marginal reduction, whereas Qwen2 degrades substantially. This suggests that composing multiple equivalence transformations introduces an effective distribution shift beyond isolated rewrites, likely through increased operator density and structural complexity despite semantic equivalence.
Implications and Diagnostic Value
The variant-wise profile provides a compact diagnostic of three failure modes: (i) missing-link sensitivity under essential rule deletion, indicating weak verification of minimal supporting premises; (ii) inconsistency handling under contradiction injection, where models fail to revise or withhold conclusions when counter-evidence is present; and (iii) compositional sensitivity, where multi-law rewrites can degrade performance even when each rewrite is semantics-preserving. The strong gap between single-law and multi-law conditions cautions against equating robustness to isolated perturbations with robustness under composition. These diagnostics motivate targeted interventions such as deletion-aware training, explicit proof/rationale verification, and stress-testing robustness under composed transformations rather than single perturbations alone.
@@ Conclusion
This work introduced a controlled framework for evaluating LLM generalisation under systematic perturbations of rule-based reasoning. By constructing multi-step inference chains and generating variants involving rule deletion, contradiction injection, and multiple families of logic-preserving transformations, we isolate behaviours that are often obscured in more heterogeneous benchmarks. Across BERT, Qwen2, and LLaMA-like models, we observe Acc=1.0000=1.0000 on the base split and no degradation under redundant rule deletion, but pronounced failures under essential rule deletion and uniform breakdown under explicit contradictions. For equivalence-based rewrites, performance is largely preserved for single-law transformations, while composed (multi-law) rewrites reveal model-dependent sensitivity. These findings underscore the gap between robustness to semantic-preserving reformulations and reliable reasoning under missing, inconsistent, or compositionally transformed evidence, and motivate methods that explicitly address premise tracking and inconsistency management.